OpenStudy (anonymous):

Which of the following is the solution to -12 + r = 3? A r = 15 B r = -15 C r = 9 D r = -9

OpenStudy (anonymous):

Isolate the variable by adding 12 to 3. -12 + 12 + r = 3 + 12 r=15

OpenStudy (anonymous):

thanx :)

OpenStudy (anonymous):

no problem:)